# GC-Bench: An Open and Unified Benchmark for Graph Condensation

### v1

@@ 1 Introduction
Data are the driving force behind advancements in machine learning, especially with the advancement of large models. However, the rapidly increasing size of datasets presents challenges in management, storage, and transmission. It also makes model training more costly and time-consuming. This issue is particularly pronounced in the graph domain, where larger datasets mean more large-scale structures, making it challenging to train models in environments with limited resources. Compared with other data types (e.g., images and texts), the irregular graph data consists of discrete nodes connected by independent edges, reflecting unique characteristics (e.g., sparsity and connectivity) in non-Euclidean space. Different from the former graph reduction works like graph coarsening, graph sparsification, and graph pooling, Graph condensation (GC) works [39, 10] are in a promising direction and show tremendous performance by condensing the large-scale dataset into a generated smaller one preserving the original datasets’ valuable information.
@@
Research Gap. Although several studies aim to comprehensively discuss existing GC methods [39, 10], they either overlook graph-specific properties or lack systematic experimentation. This discrepancy highlights a significant gap in the literature, partly due to limitations in datasets and evaluation dimensions. A concurrent work [23] analyzed the performance of node-level GC methods but it included only a subset of representative methods, lacking an analysis of graph-level methods, a deep structural analysis, and an assessment of the generalizability of the methods. To bridge this gap, we introduce GC-Bench, an open and unified benchmark to systematically evaluate existing graph condensation methods focusing on the following aspects: ❶ Effectiveness: the progress in GC, and the impact of structure and initialization on GC; ❷ Transferability: the transferability of GC methods across backbone architectures and downstream tasks; ❸ Efficiency: the time and space efficiency of GC methods. The contributions of GC-Bench are as follows:
@@ Methods
Distribution matching GCDM [20], DM [24, 22]
@@ 2.2 Benchmark Algorithms
We selected 12 representative and competitive GC methods across 6 categories for evaluation. The main ideas of these methods are shown in Figure 1. The evaluated methods include: (1) traditional core-set methods including Random, Herding [36], K-Center [30], (2) gradient matching methods including DosCond [14], GCond [15] and SGDD [42], (3) trajectory matching methods including SFGC [47] and GEOM [45], (4) distribution matching methods including GCDM [20] and DM [22, 24], (5) Kernel Ridge Regression (KRR) based method KiDD [41], and (6) Computation Tree Compression (CTC) method Mirage [11]. The details of evaluated methods are in Appendix A.2.
@@ 3.4 Transferability of Backbone Model Architectures (RQ4)
We adopt one model (SGC or Graph Transformer (GTrans)) as the backbone for condensation and use the various models in downstream tasks evaluation. Details and additional results are in Appendix B.4.
As shown in Figure LABEL:fig:arch-SGC and  LABEL:fig:arch-GTrans, each column shows the generalization performance of a condensed graph generated by different methods for various downstream models. We can observe that datasets condensed with SGC generally maintain performance when transferred across models. However, datasets condensed with Graph Transformer consistently underperform across various methods, and other models also exhibit reduced performance when adapted to Graph Transformer. Intuitively, SGC’s basic neighbor message-passing strategy may overlook global dependencies critical to more complex models, and similarly, complex models may not perform well when adapted to simpler models. As we can observe, DosCond exhibits generally better transferability compared to other gradient-matching methods. Since it can be regarded as the one-step gradient matching variant of GCond, we further test the impact of gradient matching steps on transferability (Figure LABEL:fig:Gcond-steps). Increasing the number of matching steps was found to correlate with reduced performance across architectures, indicating that extensive gradient matching may encode model-specific biases.
@@ 3.6 Efficiency and Scalability (RQ6)
In this subsection, we evaluate the condensation time and memory consumption of GC methods. The results on ogbn-arxiv are shown in Figure 5, where the x-axis denotes the overall condensation time (min) when achieving the best validation performance, the y-axis denotes the test accuracy (%), the inner size of the marker represents the peak CPU memory usage (MB), while the outer size represents the peak GPU memory usage (MB). As we can observe, the gradient matching methods have higher time and space consumption compared to other types of methods. Additionally, as shown in Table 2, the gradient matching GC methods and distribution matching GC methods result in OOM (Out of Memory) on large-scale datasets under large condensation ratios, which contradicts the original intent of graph condensation to reduce computation and storage costs for large graphs. More detailed results in Appendix B.6.
@@ A.2 Algorithms
DM [22, 24] Random Sample GNN NC ICDM, 2023
@@
      DM [22, 24]: DM can be regarded as a one-step variant of GCDM. In DM, the optimization is centered on the initial parameters. Notably, in [22] and [24], DM does not learn any structures for efficiency. However, for better comparison in our experiments, we continue to learn the adjacency matrix.

### v2

@@ 1 Introduction
Data are the driving force behind advancements in machine learning, especially with the advancement of large models. However, the rapidly increasing size of datasets presents challenges in management, storage, and transmission. It also makes model training more costly and time-consuming. This issue is particularly pronounced in the graph domain, where larger datasets mean more large-scale structures, making it challenging to train models in environments with limited resources. Compared to graph coarsening, which groups nodes into super nodes, and sparsification, which selects a subset of edges, graph condensation [39; 10] synthesizes a smaller, informative graph that retains enough data for models to perform comparably to using the full dataset.
@@
Research Gap. Although several studies aim to comprehensively discuss existing GC methods [39; 10], they either overlook graph-specific properties or lack systematic experimentation. This discrepancy highlights a significant gap in the literature, partly due to limitations in datasets and evaluation dimensions. A concurrent work [23] analyzed the performance of node-level GC methods but it included only a subset of representative methods, lacking an analysis of graph-level methods, a deep structural analysis, and an assessment of the generalizability of the methods. To bridge this gap, we introduce GC-Bench, an open and unified benchmark to systematically evaluate existing graph condensation methods focusing on the following aspects: ❶ Effectiveness: the progress in GC, and the impact of structure and initialization on GC; ❷ Transferability: the transferability of GC methods across backbone architectures and downstream tasks; ❸ Efficiency: the time and space efficiency of GC methods. The contributions of GC-Bench are as follows:
@@ Methods
Distribution matching GCDM [20], DM [24; 22]
@@ 2.2 Benchmark Algorithms
We selected 12 representative and competitive GC methods across 6 categories for evaluation. The main ideas of these methods are shown in Figure 1. The evaluated methods include: (1) traditional core-set methods including Random, Herding [36], K-Center [30], (2) gradient matching methods including DosCond [14], GCond [15] and SGDD [42], (3) trajectory matching methods including SFGC [47] and GEOM [45], (4) distribution matching methods including GCDM [20] and DM [22; 24], (5) Kernel Ridge Regression (KRR) based method KiDD [41], and (6) Computation Tree Compression (CTC) method Mirage [11]. The details of evaluated methods are in Appendix A.2.
@@ 3 Experimental Results and Analysis
*Mirage cannot directly generate graphs with the required ratio. Parameter search aligns generated graphs with DosCond disk usage (see Appendix B.1).   ‘-’ denotes results unavailable due to recursive limits reached in MP Tree search.
@@ 3.4 Transferability of Backbone Model Architectures (RQ4)
We adopt one model (SGC or Graph Transformer) as the backbone for condensation and use the various models in downstream tasks evaluation. Details and additional results are in Appendix B.4.
As shown in Figure LABEL:fig:arch-SGC and  LABEL:fig:arch-GTrans, each column shows the generalization performance of a condensed graph generated by different methods for various downstream models. We can observe that datasets condensed with SGC generally maintain performance when transferred across models. However, datasets condensed with Graph Transformer (GTrans) consistently underperform across various methods, and other models also exhibit reduced performance when adapted to Graph Transformer. Intuitively, SGC’s basic neighbor message-passing strategy may overlook global dependencies critical to more complex models, and similarly, complex models may not perform well when adapted to simpler models. As we can observe, DosCond exhibits generally better transferability compared to other gradient-matching methods. Since it can be regarded as the one-step gradient matching variant of GCond, we further test the impact of gradient matching steps on transferability (Figure LABEL:fig:Gcond-steps). Increasing the number of matching steps was found to correlate with reduced performance across architectures, indicating that extensive gradient matching may encode model-specific biases.
@@ 3.6 Efficiency and Scalability (RQ6)
In this subsection, we evaluate the condensation time and memory consumption of GC methods. The results on ogbn-arxiv are shown in Figure 5, where the x-axis denotes the overall condensation time (min) when achieving the best validation performance, the y-axis denotes the test accuracy (%), the inner size of the marker represents the peak CPU memory usage (MB), while the outer size represents the peak GPU memory usage (MB). As we can observe, the gradient matching methods have higher time and space consumption compared to other types of methods. However, Table 2 shows that current gradient and distribution matching GC methods may trigger OOM (Out of Memory) errors on large datasets with high condensation ratios, making them unsuitable for large-scale scenarios, which contradicts the goal of applying graph condensation to extremely large graphs. More detailed results in Appendix B.6.
@@ 5 Conclusion and Future Works
Acknowledgements
The corresponding author is Jianxin Li. This work is supported by the NSFC through grants No.62225202 and No.62302023, the Fundamental Research Funds for the Central Universities, CAAI-MindSpore Open Fund, developed on OpenI Community. This work is also supported in part by NSF under grants III-2106758, and POSE-2346158.
@@ A.2 Algorithms
DM [22; 24] Random Sample GNN NC ICDM, 2023
@@
      DM [22; 24]: DM can be regarded as a one-step variant of GCDM. In DM, the optimization is centered on the initial parameters. Notably, in [22] and [24], DM does not learn any structures for efficiency. However, for better comparison in our experiments, we continue to learn the adjacency matrix.